Abstract Submitted for the DFD99 Meeting of The American Physical Society

Sorting Category: 74 (Theoretical)

Many-body Effects of Matrix-Inversion in Low-Reynolds-Number Hydrodynamics KENGO ICHIKI, Graduate School of Human and Environmental Studies, Kyoto University, Japan, J.F. BRADY, Department of Chemical Engineering, California Institute of Technology — In low-Reynolds-number hydrodynamics it is known that inversion of the pair-wise mobility matrix takes into account some many-body effects. In this work the precise relation between the inverse and the many-body interactions is shown theoretically. This relation is examined in detail for a three-body problem, and show that the straightforward method of reflections may fail to converge. This limitation can be overcome by using a general iterative method, such as the generalized minimum residual method.

X Prefer Oral Session Prefer Poster Session

Kengo Ichiki

ichiki@kona.jinkan.kyoto-u.ac.jp

Graduate School of Human and Environmental Studies, Kyoto University, Japan

Date submitted: August 5, 1999 Electronic form version 1.4